please help me solve this equation for y:
y(y-4)(7y+2)=0

Answer by Gogonati(855)

Set each factor on the left side to zero:
y=0; y-4=0 => y=4; 7y+12=0 => y=-12/7
Answer:The solution set is {-12/7, 0, 4}